# Nontrivial Aharonov-Bohm effect and alternating dispersion of magnons in cone-state ferromagnetic rings

### v1

@@ I Introduction
Indeed, the spectrum of the vortex-state particle includes the gyrotropic mode with low (sub-gigahertz) frequency $\omega_{g}$, corresponding to the vortex core precession around the disc center, and a system of higher modes. These modes with the frequencies $\omega_{m,n}$ are denoted by two integers, the number of radial nodes in the out-of-plane component of the dynamical magnetization (radial mode number) $n\geq 1$ and the azimuthal mode number $m$, which determines the angular dependence on this component. These modes form doublets with the splitting $\Delta\omega_{n,m}\equiv\omega_{n,|m|}-\omega_{n,-|m|}$. Current successes in the understanding of magnon modes for the vortex-state soft-magnetic nanodiscs are partly based on past investigations of these modes for easy-plane two-dimensional (2d) models with local interactions only, see [3]. Indeed, the gyroscopic mode is present for both models, with the same sense of the vortex precessional mode, which can be described on the ground of the Thiele equation for vortex core coordinate. For both models, the frequencies of modes are determined by small parameters of the problems, which are $(a/R)^{2}$ for the 2d local model and $L/R$ for soft magnetic discs, where $R$ is the radius of the disc, $a$ is the lattice constant and $L$ is the disc thickness. The higher modes were found to form a set of doublets with given value of $n$ and $m=\pm|m|$, their mean frequencies are proportional to the square root of these small parameters, $(a/R)$ and $\sqrt{L/R}$ for both models. It was found that the interaction with the vortex core is significant for higher translation modes with $m=\pm 1$, leading to doublet splitting $\Delta\omega_{n,1}$ of the order of $\omega_{g}$. Surprisingly the ratio $\Delta\omega_{1,1}/\omega_{g}\cong 3.5$ is the universal value for these two quite different models. The dynamics of vortex core with accounting for higher modes leads to higher order Thiele equation, first proposed for a local model, see [3], and confirmed by numerical simulations for vortex state soft-magnetic dots [17, 18].
@@
Here, we show that giant doublet splitting is a generic feature related to the nontrivial topology of the vortex state and also appears due to the non-local dipolar interactions. The topology of magnetic textures gives rise to a variety of fundamentally new effects, such as skew and rainbow scattering the magnon modes in the presence of a single skyrmion in a ferromagnet [26] and a chiral magnet [24]; it also results in a finite gyro coupling force in the Thiele equation [25]. Soft magnetic dots with easy-cone magnetization possess unique topological properties, absent for other cases, either skyrmions or planar vortices. To this end, we demonstrate how the topology of a vortex gives rise to the magnon analog of the Aharonov-Bohm effect that results in giant doublet splitting. We also describe the nontrivial dependence of the mean magnon frequency on the azimuthal number. The studies were conducted analytically and numerically, with good agreement between the two approaches.
@@
Therefore, in Sec. II, we discuss the general topological features of the cone vortex state in magnetic nanodots and nanorings, which ultimately give rise to the topology-induced Aharonov-Bohm effect. In Sec. III, we present an approximate analytical model of the magnon spectrum about the cone-vortex equilibrium state, including both local and non-local interactions. In order to disentangle the giant doublet splitting in the cone state from the splitting of the $m=\pm 1$ doublet, which appears already due to the presence of a vortex core, our theory is applied to a ferromagnetic ring without a core. In Sec. IV, our model is compared with exact numerical solutions of the linearized equation of motion that governs the magnon spectrum, providing good agreement. Our approximate analytical theory describes the magnon dispersion in the cone-vortex state and directly attributes its nontrivial dependence on the azimuthal mode number $m$ to the interplay between local and non-local interactions. Finally, we demonstrate how both of these interactions contribute to the topology-induced giant doublet split. Besides providing an analytical model for magnon dynamics in (cone)-vortex-state nanodots and rings, our paper contributes to the fundamental understanding of topological effects in magnon dynamics and magnetism, in general.
II Vortex structure of the ferromagnetic ring and topology-induced Aharonov-Bohm effect.
@@
We consider a soft magnetic particle in the form of a thin ring with an inner radius $R_{\mathrm{in}}$, outer radius $R_{\mathrm{out}}$, and thickness $L$ along the normal direction $\bm{e}_{z}$. In what follows, we use the cylindrical basis $(\bm{e}_{r},\bm{e}_{\chi},\bm{e}_{z})$. For inner and outer radii much larger than the exchange length $l_{0}$ of the respective material, the magnetization direction $\bm{m}(\bm{r})=\bm{M}(\bm{r})/M_{\mathrm{s}}$ of the ring circulates around its center region to close the magnetic flux. Here, $M_{\mathrm{s}}$ is the saturation magnetization. Within the standard parametrization of the normalized magnetization $\bm{m}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)$ the vortex structure can be described as $\phi=q\chi+\pi/2,\;\theta=\theta_{0}(r)$, with $q$ being an integer. Here the function $\theta_{0}(r)$ satisfies the ordinary differential equation
@@
[…]
@@
$Q=\frac{1}{4\pi}\int\limits_{S^{2}}\mathcal{Q}\,\mathbf{d}r\mathbf{d}\chi$ (3)
@@
$Q=\frac{q}{2}\int\limits_{\theta_{0}(0)}^{\theta_{0}(\infty)}\mathrm{d}(\cos\theta_{0})=\frac{q}{2}\left[\cos\theta_{0}(0)-\cos\theta_{0}(\infty)\right].$ (6)
@@
In the original Aharonov-Bohm effectm, this total flux determines the interference between charged particles traveling along different paths around a long (formally infinite) solenoid [27]. In particular, the magnetic flux is concentrated inside the solenoid while the movement of the particles takes place outside the magnetic field region where $\bm{B}=\textrm{rot}\,\bm{A}=0$. The Hamiltonian of the system explicitly contains the vector potential and the resulting Schrödinger equation has a general solution expressed in terms of Bessel function of transcendent index $J_{|m+\alpha|}$ where $\alpha=-\Phi_{s}/\Phi_{0}$ is the phase difference acquired by particle while traveling around solenoid, with $\Phi_{0}=ch/e$ being the flux quantum. When the total flux inside the solenoid $\Phi_{s}$ is a multiple of flux quantum $\Phi_{0}$ the index of Bessel function is integer-valued and the interference is constructive (the phase shift depend only on difference between the traveled paths). Non-integer $\alpha$ means that wave function acquire additional phase factor and the interference pattern shifts.
@@
Even though magnons are not charged particles, the nontrivial topological charge of the vortex core influences the magnon excitations even when the they do not enter core region. If the total flux $\Phi$ defined by Eq. (8) is not an integer multiple of the flux quantum the wave functions of the magnons become dependent on sign of the azimuthal number $m$ and splitting for modes with different sense of gyration appears. This effect is general in the sense that cone state of a magnetic particle can be created not only by a magnetic field. In fact, the same effects are obtained in a magnet with the complex crystallographic anisotropy, including second and fourth order contributions, so that the ground state is realised by the cone state [28]. In what follows we are considering vortices with the lowest energy, $q=1$.
@@ III Theoretical description of magnon modes in the ferromagnetic ring
Here $\bar{k}$ takes the eigenvalues of the Laplace operator and determines the decrease in the demagnetizing field in free space outside the ring, for details see Appx. VI.
@@
$\bm{A}(r)=-\frac{\cos\theta_{0}}{r}\bm{e}_{\chi},$ (15)
with total flux given by Eq. (8).
@@
Before performing these calculations, some conclusions can be made from the Eq. (17) for the index of the Bessel function. In general the index $\nu$ is not-integer even for the planar vortex, and the reason is the competition of magnetostatic and exchange interactions. The limit case of exchange domination corresponds to large $m$, $m^{2}>m_{c}^{2}=L/(\bar{k}l^{2}_{0})$, $\nu=|m|$, as for local easy-plane model [3], giving the rise for growing of all the frequencies with increasing of $|m|$. For small $|m|<m_{c}$, the magnetostatic interaction dominated and $\nu\sim 1$. It is giving an anomalous dispersion, decreasing of frequencies for growing $|m|$ [6, 31]. For planar vortices, $h=0$, $\nu$ depends on $|m|$ only and the doublets with $m=\pm|m|$ are not splitted. For the cone vortex state, the situation changes drastically, $\nu(m)\neq\nu(-m)$, that leads to the doublet splitting proportional to $h$. Note that the same mathematical features, namely presence of Bessel functions with non-integer index, which depends on the sign of azimuthal quantum number of electronic states, appear in the standard AB-problem [27].
@@ IV.1 Micromagnetic modeling
To confirm our theoretical predictions, we performed micromagnetic modeling using an in-house-developed dynamic-matrix method similar to the one by Grimsditch [32] or d‘Aquino [33, 34]. For this, the linearized Landau-Lifshitz equation, which describes the temporal evolution of the magnetization dynamics for small variations $\delta\bm{m}$ around some equilibrium state $\bm{m}_{0}$, is expanded in terms of linear magnon modes $\delta\bm{m}=\sum_{\nu}\bm{m}_{j}\exp(i\omega_{j}t)+\mathrm{c.c.}$ with mode index $j$, eigenfrequencies $\omega_{j}$ and spatial mode profiles $\bm{m}_{j}(\bm{r})$. With this, the linearized equation can be written as an eigenvalue problem
@@ IV.3 Doublet split
The frequencies as a functions of external out-of-plane field $\omega(H_{0})$, showing the doublet splitting, that are obtained from both theory and simulations are presented in Fig. 3. To make the comparison more clear we introduce a correction shift $C$, representing the frequency at zero field, $C=\omega{(0)}/2\pi$. In Fig. 3 we are using $C$ as a single adjustable parameter of the theory. The calculated values of $\omega(0)$ are presented in Fig. 2, where can be seen that the difference between the frequency values obtained from numerical calculations and analytics is not more than $10\%$. It is quite a good agreement for the rough model, that reduces the nonlinear problem to the solution of ordinary differential equations allowing for an analytical solution. The magnitude of the splitting is well reproduced by the theory without any adjustable parameters, because the splitting is robust due to its topological nature.

### v3

@@ I Introduction
Spectrum of the vortex-state particle includes the gyrotropic mode with low (sub-gigahertz) frequency $\omega_{g}$, corresponding to the vortex core precession around the disc center, and a system of higher modes. These modes with the frequencies $\omega_{m,n}$ are denoted by two integers, the number of radial nodes in the out-of-plane component of the dynamical magnetization (radial mode number) $n\geq 1$ and the azimuthal mode number $m$, which determines the angular dependence on this component. These modes form doublets with the splitting $\Delta\omega_{n,m}\equiv\omega_{n,|m|}-\omega_{n,-|m|}$.
The model describing magnon dynamics on a vortex background that takes into account non-local magnetostatic dipole interaction leads to integrodifferential Landau-Lifshitz equations. In general case, this problem cannot be solved analytically. Current successes in the understanding of magnon modes for the vortex-state soft-magnetic particles are partly based on past investigations of these modes for easy-plane two-dimensional (2d) models with local interactions only, see [4]. Indeed, in case of vortex-state disk the gyroscopic mode is present for both models, with the same sense of the vortex precessional mode, which can be described on the ground of the Thiele equation for vortex core coordinate. For both models, the frequencies of modes are determined by small parameters of the problems, which are $(a/R)^{2}$ for the 2d local model and $L/R$ for soft magnetic discs, where $R$ is the radius of the disc, $a$ is the lattice constant and $L$ is the disc thickness. The higher modes were found to form a set of doublets with given value of $n$ and $m=\pm|m|$, their mean frequencies are proportional to the square root of these small parameters, $(a/R)$ and $\sqrt{L/R}$ for both models. It was found that the interaction with the vortex core is significant for higher translation modes with $m=\pm 1$, leading to doublet splitting $\Delta\omega_{n,1}$ of the order of $\omega_{g}$. Surprisingly the ratio $\Delta\omega_{1,1}/\omega_{g}\cong 3.5$ is the universal value for these two quite different models. The dynamics of vortex core with accounting for higher modes leads to higher order Thiele equation, first proposed for a local model, see [4], and confirmed by numerical simulations for vortex state soft-magnetic dots [18, 19].
@@
Here, we show that giant doublet splitting is a generic feature related to the nontrivial topology of the vortex cone state, by developing a model including both local exchange and non-local dipole interactions. The topology of magnetic textures gives rise to a variety of fundamentally new effects, such as skew and rainbow scattering the magnon modes in the presence of a single skyrmion in a ferromagnet [26] and a chiral magnet [25]; it also results in a finite gyro coupling force in the Thiele equation [27]. Soft magnetic dots with easy-cone magnetization possess unique topological properties, absent for other cases, either skyrmions or planar vortices. To this end, we demonstrate how the topology of a vortex gives rise to the magnon analog of the Aharonov-Bohm effect that results in giant doublet splitting. We also describe the nontrivial dependence of the mean magnon frequency on the azimuthal number. The studies were conducted analytically and numerically, with good agreement between the two approaches.
@@
Therefore, in Sec. II, we discuss the general topological features of the cone vortex state in magnetic nanodots and nanorings, which ultimately give rise to the topology-induced Aharonov-Bohm effect. In Sec. III, we present an approximate analytical model of the magnon spectrum about the cone-vortex equilibrium state, including both local and non-local interactions. In order to disentangle the giant doublet splitting in the cone state from the splitting of the $m=\pm 1$ doublet, which appears already due to the presence of a vortex core, our theory is applied to a ferromagnetic ring without a core. In Sec. IV, our model is compared with exact numerical solutions of the linearized equation of motion that governs the magnon spectrum, providing good agreement. Our approximate analytical theory describes the magnon dispersion in the cone-vortex state and directly attributes its nontrivial dependence on the azimuthal mode number $m$ to the interplay between local and non-local interactions. Finally, we demonstrate how both of these interactions contribute to the topology-induced giant doublet split. Besides providing an analytical model for magnon dynamics in (cone)-vortex-state nanodots and rings, our paper contributes to the fundamental understanding of topological effects in magnon dynamics and magnetism, in general. We also discuss consequences on potential applications based on cone-state vortex rings.
II Vortex structure and topology-induced Aharonov-Bohm effect.
@@
We consider a soft magnetic particle in a form of a thin ring with an inner radius $R_{\mathrm{in}}$, outer radius $R_{\mathrm{out}}$, and thickness $L$. The inner and outer radii are much larger than the exchange length $l_{0}$ of the ferromagnetic material. The magnetization vector is denoted $M(\mathbf{r})$, normalized magnetization $m(\mathbf{r})=M(\mathbf{r})/M_{s}$, where $M_{\mathrm{s}}$ is the saturation magnetization. An external magnetic field $\mathbf{H}_{0}$ is applied along the axis of the ring. In what follows, we use the cylindrical basis $(\bm{e}_{r},\bm{e}_{\chi},\bm{e}_{z})$ with the $\bm{e}_{z}$ along the axis of the particle.
The ground state of the thin magnetic particle in a shape of a disk or a ring is the vortex ground state, in which the magnetization circulates around the center region to close the magnetic flux. The vortex state was obtained as a solution of the Landau-Lifshitz equations $\partial\mathbf{M}/\partial t=\gamma[\mathbf{M}\times(\delta W/\delta\mathbf{M})]$. Here $\gamma$ is the gyromagnetic ratio, and $(\delta W/\delta\mathbf{M})$ denotes the variation of energy functional of the ferromagnet. We will discuss the specific form of the functional $W=W[\mathbf{M}]$ in the next Section. Here we just note, that only the local exchange interaction contributes to the vortex solution. The contribution of the magnetostatic dipole interaction to Landau-Lifshitz equations is proportional to $\divergence\mathbf{M}$, which for the vortex ground state is zero.
Within the standard parametrization of the normalized magnetization $\bm{m}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)$ the vortex structure can be described as $\phi=q\chi+\pi/2,\;\theta=\theta_{0}(r)$, with $q$ being an integer, see Ref. [21]. Here the function $\theta_{0}(r)$ satisfies the ordinary differential equation
@@
with the natural boundary conditions $\sin\theta_{0}(0)=0$ and $\cos\theta_{0}(\infty)=h$. Here, $h$ is the dimensionless strength of the external field: $\bm{H}_{0}=4\pi M_{s}h\bm{e}_{z}$.
@@
$Q=\frac{1}{4\pi}\int\limits_{S^{2}}\mathcal{Q}r{d}r{d}\chi$ (3)
@@
$Q=\frac{q}{2}\int\limits_{\theta_{0}(0)}^{\theta_{0}(\infty)}\sin\theta_{0}{d}\theta_{0}=\frac{q}{2}\left[\cos\theta_{0}(0)-\cos\theta_{0}(\infty)\right].$ (6)
@@
In the original Aharonov-Bohm effect, this total flux determines the interference between charged particles traveling along different paths around a long (formally infinite) solenoid [28]. In particular, the magnetic flux is concentrated inside the solenoid while the movement of the particles takes place outside the magnetic field region where $\bm{B}=\textrm{rot}\,\bm{A}=0$. The Hamiltonian of the system explicitly contains the vector potential and the resulting Schrödinger equation has a general solution expressed in terms of Bessel function of transcendent index $J_{|m+\alpha|}$ where $\alpha=-\Phi_{s}/\Phi_{0}$ is the phase difference acquired by particle while traveling around solenoid, with $\Phi_{0}=2\pi c\hbar/e$ being the flux quantum. When the total flux inside the solenoid $\Phi_{s}$ is a multiple of flux quantum $\Phi_{0}$ the index of Bessel function is integer-valued and the interference is constructive (the phase shift depend only on difference between the traveled paths). Non-integer $\alpha$ means that wave function acquire additional phase factor and the interference pattern shifts.
@@
Even though magnons are not charged particles, the nontrivial topological charge of the vortex core influences the magnon excitations even when the they do not enter core region. If the total flux $\Phi$ defined by Eq. (8) is not an integer multiple of the flux quantum the wave functions of the magnons become dependent on sign of the azimuthal number $m$ and splitting for modes with different sense of gyration appears. This effect is general in the sense that cone state of a magnetic particle can be created not only by a magnetic field. In fact, the same effects are obtained in a magnet with the complex crystallographic anisotropy, including second and fourth order contributions, so that the ground state is realised by the cone state [29].
For comparison we can mention splitting that occurs due to the magnetostatic dipole interaction, but it is not associated with the Aharonov-Bohm effect [44]. In the case of topologically induced splitting, which we consider, the effect exists even for the pure local model, that does not take into account the magnetostatic interaction.
The topological splitting of magnons is a unique property of the cone-state vortex regardless of its nature, [42]. Indeed, the total flux $\Phi$ depends on the value of magnetization far from the vortex center, $\theta_{0}(\infty)$. For the planar vortices $\theta_{0}(\infty)=\pi/2$ and the total flux is the integer multiple of the flux quantum. The same situation is with skyrmions, which are characterised by $\theta_{0}(\infty)=\pi$. In both cases $\Phi$ is an integer multiple of $\Phi_{0}$ and the Aharonov-Bohm type of splitting does not appear.
In what follows we are considering vortices with the lowest energy: only vortices with $q=1$ are stabilized by the magnetostatic dipole interaction. We expect that for higher $q$ the situation will be qualitatively the same.
The structure of magnetic vortex discussed above is basically similar for disk- and ring-shaped particles, but topological properties have peculiarities. For both cases the main topological invariant is the $\pi_{1}$-topological charge, because the vortex structure is essential for magnons to be scattered on its background. The $\pi_{2}$-topological charge is an auxiliary quantity, determined by area of part of the sphere, in which the distribution of magnetization $\mathbf{M}$ is mapped. In the disk-shaped particle this mapping covers half of the sphere for planar vortex, and the part other than half for the cone vortex. Contrary, a central part of the vortex, including the vortex core, is absent in the ring, and the topological arguments about $\pi_{2}$-topological charge are not applicable in this case, in particular, the polarization $p$ is undefined. But as we will see below, this peculiarities does not change the essence of the effect we are investigating.
Indeed, in the context of the Aharonov-Bohm effect the most fundamental point of the topological analysis is the introduction of the vector-potential, [28]. In the magnon analogue of the effect the vector-potential $\bf{A}$ induced by the vortex topology can be naturally identified from the Landau-Lifshitz equations, if they are transformed to a Schrödinger-like form. We show in the next Section that it can be done for a model that takes into account both local exchange and non-local magnetostatic interactions. We show that the vector-potential induced by the vortex topology results a giant magnon splitting in a ring-shaped magnetic particle, for which a vortex core is absent. This result corresponds exactly to the letter and spirit of the Aharonov-Bohm effect, whose goal was to show the primacy and physicality of the vector potential.
@@ III Theoretical description of magnon modes in the ferromagnetic ring
Here $\bar{k}$ takes the eigenvalues of the Laplace operator and determines the decrease in the demagnetizing field in free space outside the ring, for details see App. VI.
@@
$\bm{A}(r)=-\frac{\cos\theta_{0}}{r}\bm{e}_{\chi}.$ (15)
[…]
$B=\frac{\cos\theta_{0}}{r}\delta(r-R_{\mathrm{in}})+\frac{\sin\theta_{0}}{r}\left(\frac{d\theta_{0}}{dr}\right)\Theta(r-R_{\mathrm{in}}).$ (16)
The flux $\Phi=\int\mathbf{B}d\mathbf{s}$ has the form
$\frac{\Phi}{\Phi_{0}}=\cos\theta_{0}(R_{\mathrm{in}})+\int_{\theta_{0}(R_{\mathrm{in}})}^{\theta_{0}(\infty)}\sin\theta_{0}d\theta_{0}=\cos\theta_{0}(\infty).$ (17)
For the cone vortex state created by the external field $\cos\theta_{0}(\infty)=h$ and the final expression for the total flux is $\Phi=\Phi_{0}h$. This result differs from the one obtained for the vortex in the disk, Eq. (8). In particular, for the planar vortex in a ring-shaped particle $\Phi=0$ while for the disk case $\Phi=\Phi_{0}$. But the general properties of the vortex are the same; in both cases there is no Aharonov-Bohm effect for a planar vortex.
@@
Before performing these calculations, some conclusions can be made from the Eq. (19) for the index of the Bessel function. In general the index $\nu$ is not-integer even for the planar vortex, and the reason is the competition of magnetostatic and exchange interactions. The limit case of exchange domination corresponds to large $m$, $m^{2}>m_{c}^{2}=L/(\bar{k}l^{2}_{0})$, $\nu=|m|$, as for local easy-plane model [4], giving the rise for growing of all the frequencies with increasing of $|m|$. For small $|m|<m_{c}$, the magnetostatic interaction dominated and $\nu\sim 1$. It is giving an anomalous dispersion, decreasing of frequencies for growing $|m|$ [7, 32]. For planar vortices, $h=0$, $\nu$ depends on $|m|$ only and the doublets with $m=\pm|m|$ are not splitted. For the cone vortex state, the situation changes drastically, $\nu(m)\neq\nu(-m)$, that leads to the doublet splitting proportional to $h$. Note that the same mathematical features, namely presence of Bessel functions with non-integer index, which depends on the sign of azimuthal quantum number of electronic states, appear in the standard Aharonov-Bohm problem [28].
@@ IV.1 Micromagnetic modeling
To confirm our theoretical predictions, we performed micromagnetic modeling using an in-house-developed dynamic-matrix method similar to the one by Grimsditch [33] or d‘Aquino [34, 35] . We start with the calculation of the nonlinear ground state magnetization $\mathbf{m}_{0}$, that determines the vortex structure. For this we are solving static Landau-Lifshitz equation $\mathbf{m}_{0}\times\mathbf{H}_{\text{eff}}]=0$. Then to calculate the frequencies of the magnon eigenmodes, we are solving linearized dynamic Landau-Lifshitz equation,
$\frac{\partial{\bm{m}}}{\partial t}=-\gamma\mu_{0}\big\{\bm{m}\times\bm{H}_{\text{eff}}[\bm{m}_{0}]+\bm{m}_{0}\times\delta\bm{h}[\delta\bm{m}]\big\}.$ (20)
which describes the temporal evolution of the magnetization dynamics for small variations $\delta\bm{m}$ around equilibrium distribution $\bm{m}_{0}$. The static effective field $\bm{H}_{\text{eff}}$ contains Zeemann, exchange and dipolar interactions, while its dynamic counterpart $\delta\bm{m}$ contains only the latter two (as the external field is taken as constant in time). To proceed, Eq. (20) is expanded in terms of linear magnon modes $\delta\bm{m}=\sum_{\nu}\bm{m}_{j}\exp(i\omega_{j}t)+\mathrm{c.c.}$ with mode index $j$, eigenfrequencies $\omega_{j}$ and spatial mode profiles $\bm{m}_{j}(\bm{r})$. With this, the linearized equation can be written as an eigenvalue problem
@@ IV.3 Doublet split
The frequencies as a functions of external out-of-plane field $\omega(H_{0})$, showing the doublet splitting, that are obtained from both theory and simulations are presented in Fig. 3. To make the comparison more clear we introduce a correction shift $C$, representing the frequency at zero field, $C=\omega{(0)}/2\pi$. In Fig. 3 we are using $C$ as a single adjustable parameter of the theory. The calculated values of $\omega(0)$ are presented in Fig. 2, where can be seen that the difference between the frequency values obtained from numerical calculations and analytics is not more than $10\%$. It is quite a good agreement, taking into account the simplifications we made. The magnitude of the splitting is well reproduced by the theory without any adjustable parameters, because the splitting is robust due to its topological nature.